OpenStudy (anonymous):

Solve: 2(n - 3) + 5 = 3(n - 1) A. n = 2 B. n = -2 C. n = -1 D. n = 4

OpenStudy (anonymous):

2n-6+5=3n-3 2n-1=3n-3 2=n

OpenStudy (anonymous):

2(n - 3) + 5 = 3(n - 1) 2n-6+5=3n-3 2n-1=3n-3 -1n=-2 n=2 Same as dylanhouse but I felt the need to explain more......